# Enhancing NMR Shielding Predictions of Atoms-in-Molecules Machine Learning Models with Neighborhood-Informed Representations

### v1

@@ I Introduction
Nuclear magnetic resonance (NMR) spectroscopy is a versatile, non-destructive technique for elucidating the three-dimensional structures of molecules and crystals. Ab initio NMR calculations Helgaker, Jaszuński, and Ruud (1999); Mulder and Filatov (2010), typically based on density functional theory (DFT), complement experiments by validating proposed molecular structures Nicolaou and Snyder (2005); Suyama, Gerwick, and McPhail (2011); Maier (2009), assigning chemical shifts Grimblat and Sarotti (2016), resolving cases where experimental spectra are insufficient Das, Edison, and Merz Jr (2020); Cuadrado, Cen-Pacheco, and Daranas (2024), and determining stereochemistry in diastereomers Smith and Goodman (2009, 2010). However, the broad conformational diversity of large molecules often limits the scope of such ab initio approaches, as the cost of accurate geometry optimizations and NMR parameter calculations is prohibitive for large-scale, high-throughput screening required in areas such as drug discovery and metabolite annotation.
@@
In contrast to DFT-based modeling, empirical methods such as those in ChemDraw Herlt (2025) and MarvinSketch Csizmadia (1999) predict NMR chemical shifts directly from atomic connectivities, without requiring three-dimensional structures. ChemDraw employs an additive model with thousands of parameters fitted to the curated ChemNMR dataset Herlt (2025), yielding typical prediction errors of 3.8 ppm for 13C shifts across several environments and 0.2–0.3 ppm for 1H shifts of CHxx groups (with C connected to 1–3 H atoms) Cheeseman and Frisch (2000). MarvinSketch relies on the hierarchical ordered spherical description of environment (HOSE) Bremser (1978), which is trained on the NMRShiftDB database Steinbeck and Kuhn (2004). While such empirical approaches offer near-instantaneous predictions and can distinguish many common isomers, they frequently break down for under-represented or complex environments that occur across the vast chemical space, such as those enumerated in the GDB17 database Ruddigkeit et al. (2012) of small organic molecules with up to 17 CONF atoms.
@@
Machine learning (ML) provides a powerful data-driven alternative for modeling molecular properties Rupp et al. (2012). Once trained on first-principles calculations or other reference data, ML models deliver predictions with near-empirical speed while retaining the accuracy of the underlying target method Ramakrishnan and von Lilienfeld (2017). A growing body of work Rupp, Ramakrishnan, and Von Lilienfeld (2015); Gupta, Chakraborty, and Ramakrishnan (2021); Zou et al. (2023); Shiota, Ishihara, and Mizukami (2024); Xu et al. (2025); Ramos, Mueller, and Beran (2025); Gaumard et al. (2022); Ondar, Polynski, and Ananikov (2023); El-Samman and De Baerdemacker (2025); Paruzzo et al. (2018); Gerrard et al. (2020); Büning, Grimme, and Bursch (2024); Howarth, Ermanis, and Goodman (2020); Wang et al. (2025); Lemm, von Rudorff, and Von Lilienfeld (2024); Cordova et al. (2022, 2023); Unzueta and Beran (2025); Steinbeck and Kuhn (2004); Yuan et al. (2025) has demonstrated the applicability of ML combined with quantum chemistry for predicting a wide range of NMR properties in molecules and solids. To support such models, several dedicated datasets have been developed and curated, including QM9NMR Gupta, Chakraborty, and Ramakrishnan (2021), IMPRESSION Gerrard et al. (2020), NMRShiftDB Steinbeck and Kuhn (2004), and QMe14S Yuan et al. (2025), which provide thousands of examples required to reduce training set bias. Among these, QM9NMR Gupta, Chakraborty, and Ramakrishnan (2021) is one of the largest, containing structures and NMR shielding parameters for 1H, 13C, 15N, 17O, and 19F in 130,831 molecules from the QM9 dataset Ramakrishnan et al. (2014) of small organic molecules with up to 9 CONF atoms, including 831,925 13C shieldings. This dataset has enabled the training of ML models with up to 100,000 (i.e., 100k) examples. Gupta et al. Gupta, Chakraborty, and Ramakrishnan (2021) trained a kernel ridge regression (KRR) model on 100k QM9NMR entries using the Faber–Christensen–Huang–Lilienfeld (FCHL) many-body descriptor Faber et al. (2018), achieving an out-of-sample error of 1.88 ppm for 13C shifts of 50k hold-out atoms. Using the same split, Shiota et al. Shiota, Ishihara, and Mizukami (2024) reported a similar error of 1.87 ppm with a KRR model based on a graph neural network–derived transfer learning embedding and the MACE-OFF23-small descriptor (the “small” variant of a transferable organic force field derived from the message passing atomic cluster expansion architecture Kovács et al. (2025)). Notably, when applied to drug molecules outside the training domain, the model of Shiota et al. delivered smaller errors than that of Gupta et al.
@@
The objective of this study is to develop systematically improved atom-centered structural descriptors for predicting NMR chemical shifts, trained on the QM9NMR dataset. We introduce atomic variants of the Coulomb matrix (CM) Rupp et al. (2012) and bag-of-bonds (BoB) Hansen et al. (2015) descriptors in both discrete and continuous formulations. To further capture local chemical environments, we augment these atomic descriptors with information on nearest neighbor atoms. The resulting representations achieve lower out-of-sample errors than existing approaches, establishing new benchmarks for diverse classes of organic molecules. The remainder of this article is organized as follows. Section II introduces the datasets used in this study and outlines the computational details. Section III describes the kernel ridge regression (KRR) approach and the construction of the atomic descriptors. Section IV presents the optimization of hyperparameters, evaluates the efficiency of the descriptors, and discusses model performance for predicting 13C chemical shifts across different datasets.
@@ II Datasets and Computational Details
We collected QM9NMR, Drug12, Drug40, and GDBm datasets from the QM9NMR dataset’s repositoryqm9 (2021) developed in a previous studyGupta, Chakraborty, and Ramakrishnan (2021). QM9 molecules were optimized using the B3LYP/6-31G(2df,p) DFT/basis set combination. For ML modeling, the target quantities used are NMR parameters of QM9 molecules, calculated at the mPW1PW91/6-311+G(2d,p) level, using geometries in QM9 determined with the B3LYP/6-31G(2df,p) method. These values were collected from the QM9NMR datasetqm9 (2021). In this study, we trained ML models using descriptors based on minimum energy geometries calculated with the PM7 semi-empirical model, as a previous study has shown the efficiency of using descriptors derived from geometries determined with inexpensive methodsRamakrishnan et al. (2015). For QM9NMR, Drug12, Drug40, and GDBm datasets, PM7 geometries were collected from the QM9NMR datasetqm9 (2021). Table 1 summarizes the molecular datasets explored in this study, with specific details for each dataset provided in the subsections below.
@@
where tetramethylsilane (TMS) is used as the standard reference for 13C NMR. The gas-phase $\sigma_{\rm iso}(^{13}{\rm C})$ of TMS, calculated at the mPW1PW91/6-311+G(2d,p) level using the minimum-energy geometry optimized with B3LYP/6-31G(2df,p), was taken from the QM9NMR repository and used as the reference value, $\sigma_{\rm iso}^{\rm TMS}=186.9704$ ppm qm9 (2021).
@@ II.1 QM9 and QM9NMR
The QM9 dataset Ramakrishnan et al. (2014); Ramakrishnan (2024) is a widely used benchmark for ML modeling of molecular properties. Building on this resource, the QM9NMR dataset Gupta, Chakraborty, and Ramakrishnan (2021) provides isotropic NMR chemical shielding values ($\sigma_{\rm iso}$) for all nuclei present in QM9: 1,208,486 1H, 831,925 13C, 132,498 15N, 183,265 17O, and 3,036 19F. Shieldings are reported for both gas-phase molecules and in implicit-solvent environments (CCl4, THF, acetone, methanol, and DMSO).
@@
In Figure 1, the $\delta(^{13}{\rm C})$ distribution of the QM9NMR dataset is shown by carbon hybridization. The $sp^{3}$ carbons ($\sim$600k atoms) span a broad range of 0–100 ppm, while the $sp^{2}$ carbons ($\sim$177k atoms) extend over 100–225 ppm. The $sp$ carbons ($\sim$56k atoms) fall in the mid-range, 50-125 ppm, reflecting anisotropic fields associated with triple bonds. Their distribution is bimodal, with peaks corresponding to carbons in -C#N and -C#C- groups (where # denotes a triple bond in SMILES notation Weininger (1988)). The right-most peak arises from the deshielded carbon bound to electronegative nitrogen (-C#N), whereas the left-most peak corresponds to the more shielded carbon bound to another carbon (-C#C-). Because -C#N can extend only in one direction, while -C#C- can extend from both ends, the latter shows a broader distribution.
@@ II.2 Drug12 and Drug40
For validation of the ML models, two completely out-of-sample datasets, Drug12 (twelve drug molecules from Ref. Corey, Czakó, and Kürti, 2007) and Drug40 (a subset of forty molecules from the GDB17 dataset Ref. Ruddigkeit et al., 2012), are taken into consideration. These two datasets are also used in previous works Refs. Gupta, Chakraborty, and Ramakrishnan, 2021; Shiota, Ishihara, and Mizukami, 2024. In the Drug40 dataset, molecules contain 7 to 17 heavy atoms (Table 1), whereas the training of the ML model is done with at most 9 atoms in the QM9 dataset. In Drug12, the heavy atoms present in a molecule are much higher, 17 to 23. As these datasets contain a larger number of heavy atoms than the QM9 dataset, they are good candidates for the validation of the ML models.
Also, as the heavy atoms are much higher in number in these datasets, they will ensure the transferability of the ML models to large molecular compounds. That is, even though the ML model is trained on a dataset containing a low number of heavy atoms, they are transferable to larger molecules. The PM7 geometries of the drug molecules are obtained from the QM9NMR repository qm9 (2021).
@@ II.3 GDBm
As a validation set, we also employ the GDBm dataset as a validation set, where $m$ ranges from 10 to 17, denoting the number of heavy atoms in a molecule. For example, GDB13 contains molecules with up to 13 CONF atoms. For each $m$, 25 molecules are selected, yielding a total of 200 molecules in the GDBm set. This dataset was previously used as a validation benchmark in the QM9NMR study Gupta, Chakraborty, and Ramakrishnan (2021). In the present work, GDBm serves to assess the transferability and accuracy of the proposed descriptors. The PM7-optimized geometries of the GDBm molecules are obtained from the QM9NMR repository qm9 (2021).
@@ II.4 Pyrimidinone
For further validation of the ML models of $\delta(^{13}{\rm C})$ presented in this study, we consider a set of substituted pyrimidinone molecules. The dataset comprises 208 biomolecules with substitutions (-CH3, -NH2, and -F), including 16 uracil [pyrimidine-2,4(1H,3H)-dione], 64 pyrimidine-2(1H)-one, 64 pyrimidine-4(1H)-one, and 64 pyrimidine-4(3H)-one derivatives. These molecules contain 7–10 heavy atoms, making them suitable candidates for validation. This set was previously used for evaluating ML predictions of core electron binding energies Tripathy et al. (2024a). In the present work, we extend their use to assess NMR chemical shifts, thereby broadening the scope of properties available for these biologically relevant molecules.
For the pyrimidinone dataset, geometries calculated with the universal force field (UFF) were collected from a repositoryTripathy et al. (2024b) developed for a previous studyTripathy et al. (2024a). We optimized these geometries with B3LYP/6–31G(2df,p) and calculated NMR shielding parameters as analytical second derivatives of total energy with respect to magnetic moments and magnetic field Gauss (2000) with mPW1PW91/6-311+G(2d,p) using Gaussian16 Frisch et al. (2016). ML predictions of the chemical shifts of pyrimidinone molecules were done using geometries determined with PM7 using the MOPAC software Stewart (1990).
@@ III.1 Atomic representations of local chemical environments
In ML modeling of molecules and materials, a descriptor (or representation) provides a mathematical mapping of atomistic structural features. Global descriptors encode the entire molecular structure to predict properties such as total energy Hansen et al. (2015), whereas local (atomic) descriptors capture the structural environment around individual atoms to predict atom-specific properties. The latter approach is commonly referred to as atoms-in-molecules (AIM) machine learning Rupp, Ramakrishnan, and Von Lilienfeld (2015), which is particularly relevant for NMR shielding modeling. Since NMR shielding is a property of a quasi-atom within a molecule, with the observed chemical shift arising from its local chemical environment, AIM-ML offers a natural framework for such predictions. More recently, deep learning-based interatomic potential descriptors have been developed to encode atomic environments using graph neural networks (GNNs) Scarselli et al. (2008). For example, the ‘AtmEnV’ descriptor is derived from the SchNet architecture Schütt et al. (2018) to represent the local environment of an atom Tripathy et al. (2024a). Shiota et al. Shiota, Ishihara, and Mizukami (2024) employed GNN-based descriptors for chemical shift prediction, including MACE Batatia et al. (2022), which is a physics-inspired architecture. In the present study, we focus on simpler structure-based descriptors explicitly derived from ab initio geometries.
@@
In this study, we investigate the accuracy of descriptors by focusing on two key aspects: the transformation from discrete to continuous representations and the role of many-body effects captured through the nearest-neighbor chemical environment. Continuous descriptors are of constant size Collins et al. (2018) and are typically represented on radial or angular grids Becker (1977) that encode the atomistic surroundings of a query atom. In contrast, discrete descriptors are counted entity-wise (atomwise, bondwise, or via higher-order motifs), making them chemically intuitive but variable in size across molecules. For molecular applications, continuous descriptors offer a physically grounded representation of local environments, while discrete descriptors emphasize interpretable structural features that can be directly compared across systems. Despite their importance, a systematic evaluation of discrete versus continuous formulations in chemical property prediction remains lacking. Beyond chemistry, discrete representations have been widely explored. For example, human language and image processing provide canonical cases where information is encoded as a discrete set of symbols or pixels, respectively Cartuyvels, Spinks, and Moens (2021). Meyer et al. recently showed that large discrete representations can improve performance in continuous reinforcement learning Meyer, White, and Machado (2025). Descriptor performance can also be improved by incorporating higher-order many-body interaction terms, which enhance both accuracy and size efficiency Huang, Symonds, and von Lilienfeld (2020). A well-known example is the spectrum of London Axilrod–Teller–Muto (SLATM) descriptor Huang and von Lilienfeld (2020), which includes the three-body Axilrod–Teller–Muto (ATM) potential. Similarly, the bonds and angles ML (BAML) descriptor Huang and Von Lilienfeld (2016) encodes atom, bond, angle, and torsional terms derived from the many-body expansions of the universal force field (UFF).
We revisit the atomic Coulomb matrix (aCM) Rupp et al. (2012), introduce the atomic variant of bag-of-bonds (BoB) Hansen et al. (2015), denoted aBoB, and develop their continuous counterparts as illustrated in Figure 2. Both aCM and aBoB are based on two-body Coulombic interactions, whereas the atomic SLATM (aSLATM) Huang and von Lilienfeld (2020) additionally incorporates three-body Axilrod–Teller–Muto (ATM) terms. We freshly implement aSLATM by integrating the three-body contribution in the $\cos\theta$ domain rather than the $\theta$ domain, leading to a significant reduction in descriptor size. To enhance representability, we further augment each atomic descriptor with contributions from neighboring atomic centers.

### v3

@@ I Introduction
Nuclear magnetic resonance (NMR) spectroscopy is a versatile, non-destructive technique for elucidating the three-dimensional structures of molecules and crystals. Ab initio NMR calculations helgaker1999ab; mulder2010nmr, typically based on density functional theory (DFT), complement experiments by validating proposed molecular structures nicolaou2005chasing; suyama2011survey; maier2009structural, assigning chemical shifts grimblat2016computational, resolving cases where experimental spectra are insufficient das2020metabolite; cuadrado2024computationally, and determining stereochemistry in diastereomers smith2009assigning; smith2010assigning. However, the broad conformational diversity of large molecules often limits the scope of such ab initio approaches, as the cost of accurate geometry optimizations and NMR parameter calculations is prohibitive for large-scale, high-throughput screening required in areas such as drug discovery and metabolite annotation.
@@
In contrast to DFT-based modeling, empirical methods such as those in ChemDraw chemnmr and MarvinSketch csizmadia1999marvinsketch predict NMR chemical shifts directly from atomic connectivities, without requiring three-dimensional structures. ChemDraw employs an additive model with thousands of parameters fitted to the curated ChemNMR dataset chemnmr, yielding typical prediction errors of 3.8 ppm for 13C shifts across several environments and 0.2–0.3 ppm for 1H shifts of CHxx groups (with C connected to 1–3 H atoms) cheeseman2000predicting. MarvinSketch relies on the hierarchical ordered spherical description of environment (HOSE) bremser1978hose, which is trained on the NMRShiftDB database steinbeck2004nmrshiftdb. While such empirical approaches offer near-instantaneous predictions and can distinguish many common isomers, they frequently break down for under-represented or complex environments that occur across the vast chemical space, such as those enumerated in the GDB17 database ruddigkeit2012enumeration of small organic molecules with up to 17 CONF atoms.
@@
Machine learning (ML) provides a powerful data-driven alternative for modeling molecular properties rupp2012fast. Once trained on first-principles calculations or other reference data, ML models deliver predictions with near-empirical speed while retaining the accuracy of the underlying target method ramakrishnan2017machine. A growing body of work rupp2015machine; gupta2021revving; zou2023deep; shiota2024universal; xu2025toward; ramos2025interplay; gaumard2022regression; ondar2023predicting; el2025amide; paruzzo2018chemical; gerrard2020impression; buning2024machine; howarth2020dp4; wang2025nmrextractor; lemm2024impact; cordova2022machine; cordova2023atomic; unzueta2025predicting; steinbeck2004nmrshiftdb; yuan2025qme14s has demonstrated the applicability of ML combined with quantum chemistry for predicting a wide range of NMR properties in molecules and solids. To support such models, several dedicated datasets have been developed and curated, including QM9NMR gupta2021revving, IMPRESSION gerrard2020impression, NMRShiftDB steinbeck2004nmrshiftdb, and QMe14S yuan2025qme14s, which provide thousands of examples required to reduce training set bias. Among these, QM9NMR gupta2021revving is one of the largest, containing structures and NMR shielding parameters for 1H, 13C, 15N, 17O, and 19F in 130,831 molecules from the QM9 dataset ramakrishnan2014quantum of small organic molecules with up to 9 CONF atoms, including 831,925 13C shieldings. This dataset has enabled the training of ML models with up to 100,000 (i.e., 100k) examples. Gupta et al. gupta2021revving trained a kernel ridge regression (KRR) model on 100k QM9NMR entries using the Faber–Christensen–Huang–Lilienfeld (FCHL) many-body descriptor faber2018alchemical, achieving an out-of-sample error of 1.88 ppm for 13C shifts of 50k hold-out atoms. Using the same split, Shiota et al. shiota2024universal reported a similar error of 1.87 ppm with a KRR model based on a graph neural network–derived transfer learning embedding and the MACE-OFF23-small descriptor (the “small” variant of a transferable organic force field derived from the message passing atomic cluster expansion architecture kovacs2025mace). Notably, when applied to drug molecules outside the training domain, the model of Shiota et al. delivered smaller errors than that of Gupta et al.
@@
The objective of this study is to develop systematically improved atom-centered structural descriptors for predicting NMR chemical shifts, trained on the QM9NMR dataset. We introduce atomic variants of the Coulomb matrix (CM) rupp2012fast and bag-of-bonds (BoB) hansen2015machine descriptors in both discrete and continuous formulations. To further capture local chemical environments, we augment these atomic descriptors with information on nearest neighbor atoms. The resulting representations achieve lower out-of-sample errors than existing approaches, establishing new benchmarks for diverse classes of organic molecules. The remainder of this article is organized as follows. Section II introduces the datasets used in this study and outlines the computational details. Section III describes the kernel ridge regression (KRR) approach and the construction of the atomic descriptors. Section IV presents the optimization of hyperparameters, evaluates the efficiency of the descriptors, and discusses model performance for predicting 13C chemical shifts across different datasets.
@@ II Datasets and Computational Details
We collected QM9NMR, Drug12, Drug40, and GDBm datasets from the QM9NMR dataset’s repositoryqm9nmr developed in a previous studygupta2021revving. QM9 molecules were optimized using the B3LYP/6-31G(2df,p) DFT/basis set combination. For ML modeling, the target quantities used are NMR parameters of QM9 molecules, calculated at the mPW1PW91/6-311+G(2d,p) level, using geometries in QM9 determined with the B3LYP/6-31G(2df,p) method. These values were collected from the QM9NMR datasetqm9nmr. In this study, we trained ML models using descriptors based on minimum energy geometries calculated with the PM7 semi-empirical model, as a previous study has shown the efficiency of using descriptors derived from geometries determined with inexpensive methodsramakrishnan2015big. For QM9NMR, Drug12, Drug40, and GDBm datasets, PM7 geometries were collected from the QM9NMR datasetqm9nmr. We computed NMR parameters for 208 biologically relevant pyrimidinone molecules using the same protocol employed in the QM9NMR dataset gupta2021revving. Specifically, NMR properties were calculated at the mPW1PW91/6-311+G(2d,p) level of theory, using minimum-energy geometries obtained with B3LYP/6-31G(2df,p) from the QM9 dataset ramakrishnan2014quantum. Table 1 summarizes the molecular datasets explored in this study, with specific details for each dataset provided in the subsections below.
@@
where tetramethylsilane (TMS) is used as the standard reference for 13C NMR. The gas-phase $\sigma_{\rm iso}(^{13}{\rm C})$ of TMS, calculated at the mPW1PW91/6-311+G(2d,p) level using the minimum-energy geometry optimized with B3LYP/6-31G(2df,p), was taken from the QM9NMR repository and used as the reference value, $\sigma_{\rm iso}^{\rm TMS}=186.9704$ ppm qm9nmr.
For the $\Delta$-ML modeling ramakrishnan2015big, baseline NMR calculations were carried out at the HF/STO-3G level using PM7-optimized geometries obtained from the QM9NMR repository qm9nmr. The HF baseline was originally chosen because it is conceptually simple, implementation-independent compared to DFT methods, and widely available across quantum chemistry software packages. To further explore alternatives, we also evaluated the cost of using the mPW1PW91 functional with the minimal basis set, STO-3G for a representative molecule containing 9 CONF atoms (SMILES: CC1CCCCC1C=O), NMR shielding tensor calculations at the mPW1PW91/6-311+G(2d,p) and mPW1PW91/STO-3G levels require 776.5 and 21.6 seconds of CPU time, respectively, amounting to a speedup of roughly 35-fold. This substantial reduction in computational cost highlights the suitability of minimal-basis calculations as baselines for $\Delta$-ML modeling of NMR chemical shifts.
Baseline NMR calculations for all QM9 molecules were newly performed at the mPW1PW91/STO-3G level using Gaussian16 frisch2016gaussian and the same PM7-level geometries obtained from the QM9NMR repository qm9nmr. Since mPW1PW91 is also the functional used to generate the target NMR properties, the resulting $\Delta$-ML model learns to correct both (i) the geometric difference between PM7 and B3LYP/6-31G(2df,p) structures and (ii) the basis-set change from STO-3G to 6-311+G(2d,p).
@@ II.1 QM9 and QM9NMR
The QM9 dataset ramakrishnan2014quantum; qm9pack is a widely used benchmark for ML modeling of molecular properties. Building on this resource, the QM9NMR dataset gupta2021revving provides isotropic NMR chemical shielding values ($\sigma_{\rm iso}$) for all nuclei present in QM9: 1,208,486 1H, 831,925 13C, 132,498 15N, 183,265 17O, and 3,036 19F. Shieldings are reported for both gas-phase molecules and in implicit-solvent environments (CCl4, THF, acetone, methanol, and DMSO).
@@
In Figure 1, the $\delta(^{13}{\rm C})$ distribution of the QM9NMR dataset is shown by carbon hybridization. The $sp^{3}$ carbons ($\sim$600k atoms) span a broad range of 0–100 ppm, while the $sp^{2}$ carbons ($\sim$177k atoms) extend over 100–225 ppm. The $sp$ carbons ($\sim$56k atoms) fall in the mid-range, 50-125 ppm, reflecting anisotropic fields associated with triple bonds. Their distribution is bimodal, with peaks corresponding to carbons in -C#N and -C#C- groups (where # denotes a triple bond in SMILES notation weininger1988smiles). The right-most peak arises from the deshielded carbon bound to electronegative nitrogen (-C#N), whereas the left-most peak corresponds to the more shielded carbon bound to another carbon (-C#C-). Because -C#N can extend only in one direction, while -C#C- can extend from both ends, the latter shows a broader distribution.
@@ II.2 Drug12 and Drug40
For validation of the ML models, two completely out-of-sample datasets, Drug12 (twelve drug molecules from Ref. corey2007molecules) and Drug40 (a subset of forty molecules from the GDB17 dataset Ref. ruddigkeit2012enumeration), are taken into consideration. These two datasets are also used in previous works Refs. gupta2021revving; shiota2024universal. In the Drug40 dataset, molecules contain 7 to 17 heavy atoms (Table 1), whereas the training of the ML model is done with at most 9 atoms in the QM9 dataset. In Drug12, the heavy atoms present in a molecule are much higher, 17 to 23. As these datasets contain a larger number of heavy atoms than the QM9 dataset, they are good candidates for the validation of the ML models.
Also, as the heavy atoms are much higher in number in these datasets, they will ensure the transferability of the ML models to large molecular compounds. That is, even though the ML model is trained on a dataset containing a low number of heavy atoms, they are transferable to larger molecules. The PM7 geometries of the drug molecules are obtained from the QM9NMR repository qm9nmr.
@@ II.3 GDBm
As a validation set, we also employ the GDBm dataset as a validation set, where $m$ ranges from 10 to 17, denoting the number of heavy atoms in a molecule. For example, GDB13 contains molecules with up to 13 CONF atoms. For each $m$, 25 molecules are selected, yielding a total of 200 molecules in the GDBm set. This dataset was previously used as a validation benchmark in the QM9NMR study gupta2021revving. In the present work, GDBm serves to assess the transferability and accuracy of the proposed descriptors. The PM7-optimized geometries of the GDBm molecules are obtained from the QM9NMR repository qm9nmr.
@@ II.4 Pyrimidinone
For further validation of the ML models of $\delta(^{13}{\rm C})$ presented in this study, we consider a set of substituted pyrimidinone molecules. The dataset comprises 208 biomolecules with substitutions (-CH3, -NH2, and -F), including 16 uracil [pyrimidine-2,4(1H,3H)-dione], 64 pyrimidine-2(1H)-one, 64 pyrimidine-4(1H)-one, and 64 pyrimidine-4(3H)-one derivatives. These molecules contain 7–10 heavy atoms, making them suitable candidates for validation. This set was previously used for evaluating ML predictions of core electron binding energies tripathy2024chemical. In the present work, we extend their use to assess NMR chemical shifts, thereby broadening the scope of properties available for these biologically relevant molecules.
For the pyrimidinone dataset, geometries calculated with the universal force field (UFF) were collected from a repositorycebeconf developed for a previous studytripathy2024chemical. We optimized these geometries with B3LYP/6–31G(2df,p) and calculated NMR shielding parameters as analytical second derivatives of total energy with respect to magnetic moments and magnetic field gauss2000molecular with mPW1PW91/6-311+G(2d,p) using Gaussian16 frisch2016gaussian. ML predictions of the chemical shifts of pyrimidinone molecules were done using geometries determined with PM7 using the MOPAC software stewart1990mopac.
@@ III.1 Atomic representations of local chemical environments
In ML modeling of molecules and materials, a descriptor (or representation) provides a mathematical mapping of atomistic structural features. Global descriptors encode the entire molecular structure to predict properties such as total energy hansen2015machine, whereas local (atomic) descriptors capture the structural environment around individual atoms to predict atom-specific properties. The latter approach is commonly referred to as atoms-in-molecules (AIM) machine learning rupp2015machine, which is particularly relevant for NMR shielding modeling. Since NMR shielding is a property of a quasi-atom within a molecule, with the observed chemical shift arising from its local chemical environment, AIM-ML offers a natural framework for such predictions. More recently, deep learning-based interatomic potential descriptors have been developed to encode atomic environments using graph neural networks (GNNs) scarselli2008graph. For example, the ‘AtmEnV’ descriptor is derived from the SchNet architecture schutt2018schnet to represent the local environment of an atom tripathy2024chemical. Shiota et al. shiota2024universal employed GNN-based descriptors for chemical shift prediction, including MACE batatia2022mace, which is a physics-inspired architecture. In the present study, we focus on simpler structure-based descriptors explicitly derived from ab initio geometries.
@@
In this study, we investigate the accuracy of descriptors by focusing on two key aspects: the transformation from discrete to continuous representations and the role of many-body effects captured through the nearest-neighbor chemical environment. Continuous descriptors are of constant size collins2018constant and are typically represented on radial or angular grids becker1977synthesis that encode the atomistic surroundings of a query atom. In contrast, discrete descriptors are counted entity-wise (atomwise, bondwise, or via higher-order motifs), making them chemically intuitive but variable in size across molecules. For molecular applications, continuous descriptors offer a physically grounded representation of local environments, while discrete descriptors emphasize interpretable structural features that can be directly compared across systems. Despite their importance, a systematic evaluation of discrete versus continuous formulations in chemical property prediction remains lacking. Beyond chemistry, discrete representations have been widely explored. For example, human language and image processing provide canonical cases where information is encoded as a discrete set of symbols or pixels, respectively cartuyvels2021discrete. Meyer et al. recently showed that large discrete representations can improve performance in continuous reinforcement learning meyer2025harnessing. Descriptor performance can also be improved by incorporating higher-order many-body interaction terms, which enhance both accuracy and size efficiency huang2020quantummachine. A well-known example is the spectrum of London Axilrod–Teller–Muto (SLATM) descriptor huang2020quantum, which includes the three-body Axilrod–Teller–Muto (ATM) potential. Similarly, the bonds and angles ML (BAML) descriptor huang2016communication encodes atom, bond, angle, and torsional terms derived from the many-body expansions of the universal force field (UFF).
We revisit the atomic Coulomb matrix (aCM) rupp2012fast, introduce the atomic variant of bag-of-bonds (BoB) hansen2015machine, denoted aBoB, and develop their continuous counterparts as illustrated in Figure 2. Both aCM and aBoB are based on two-body Coulombic interactions, whereas the atomic SLATM (aSLATM) huang2020quantum additionally incorporates three-body Axilrod–Teller–Muto (ATM) terms. We freshly implement aSLATM by integrating the three-body contribution in the $\cos\theta$ domain rather than the $\theta$ domain, leading to a significant reduction in descriptor size. To enhance representability, we further augment each atomic descriptor with contributions from neighboring atomic centers.